# Computing Anharmonic Infrared Spectra of Polycyclic Aromatic Hydrocarbons Using Machine-Learning Molecular Dynamics

### v1

@@ IV Conclusions
Despite the impressive efficiency of MLMD, several limitations persist. First, the current model was trained exclusively on neutral PAHs with 12C and 13N and 1H and does not account for charged molecules or isotopologues, both of which play a significant role in astronomical contexts. This limitation stems from the scarcity of data for the IR spectra of charged PAHs and isotopologues, as well as challenges in incorporating charge state and isotope information into molecular descriptors. Second, because of the data-driven nature of the approach, the model may exhibit increased uncertainty when predicting spectra for molecules that deviate substantially from those in the training dataset. Moreover, training data calculated using DFT at the B3LYP/4-31G level may introduce uncertainty in predicting spectra for molecules containing heteroatoms, as the polarization effects are not adequately captured by this relatively simple basis set. We are actively working to address these challenges, with the goal of further enhancing MLMD’s capabilities for IR spectrum computation.
@@
The source code and the MLMD model to calculate the anharmonic spectra of PAHs using MLMD are freely available at Zenodo. The model is trained and ready for use without the need for a further ML procedure. It will be continuously updated to improve predictive performance. This repository also includes the spectral data for the 1 704 theoretically-calculated and 49 experimentally-tested PAHs.

### v2

@@ IV Conclusions
Despite the impressive efficiency of MLMD, several limitations persist. First, the current model was trained exclusively on neutral PAHs with natural chemical elements and does not account for charged molecules or isotopologues, both of which may play a significant role in astronomical contexts. This limitation stems from the scarcity of data for the IR spectra, as well as challenges in incorporating charge state and isotope information into molecular descriptors. Second, because of the data-driven nature of the approach, the model may exhibit increased uncertainty when predicting spectra for molecules that deviate substantially from those in the training dataset. We are actively working to address these challenges, with the goal of further enhancing MLMD’s capabilities for IR spectrum computation.
@@
The source code and the MLMD model to calculate the anharmonic spectra of PAHs using MLMD are freely available as the supplementary information of this article. The model is trained and ready for use without the need for a further ML procedure. It will be continuously updated to improve predictive performance. The supplementary file also includes the spectral data for the 1 704 theoretically-calculated and 49 experimentally-tested PAHs.